OpenStudy (kpop4life123):

Which set of line segments could create a right triangle? (4 points) 5, 6, 11 5, 9, 10 5, 13, 18 5, 12, 13

Directrix (directrix):

Use this: Pythagorean Thereom converse: If the square of one side of a triangle is equal to the sum of the squares of the other two sides, then the triangle is a right triangle.

OpenStudy (kpop4life123):

i dont understand lol

Directrix (directrix):

Does 13^2 = 12^2 + 5^2

Directrix (directrix):

@Kpop4life123 Am waiting for you to crank out 13^2 = 12^2 = and 5^2=

OpenStudy (kpop4life123):

13^2=169 12^2=144 5^2=25

OpenStudy (kpop4life123):

@Directrix

OpenStudy (kpop4life123):

???????????????????

OpenStudy (kpop4life123):

@goalieboy

Directrix (directrix):

13^2=169 12^2=144 5^2=25 Does 169 = 144 + 25 ? If so, the Converse of the Pythagorean Theorem applies and the lengths are lengths of the sides of a right triangle.

Directrix (directrix):

@Kpop4life123

Directrix (directrix):

Add those numbers and let's get a wrap on this problem.

OpenStudy (goalieboy):

what do you need help with